# Indecomposable involutive set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v2

@@ Abstract.
The quantum Yang–Baxter equation is a braiding condition on vector spaces which is of high relevance in several fields of mathematics, such as knot theory and quantum group theory. Their combinatorial counterpart are set-theoretic solutions to the Yang–Baxter equation, whose investigation is strongly driven by the study of algebraic objects called (skew) braces. In this article, we focus on indecomposable involutive non-degenerate set-theoretic solutions to the Yang–Baxter equation. More specifically, through a thorough analysis of their associated braces, we give a full classification of those which are of size $p^{2}$, for $p$ a prime.
@@ Introduction
Outside of physics, solutions to this equation have been recognized to be of importance in low-dimensional topology, as results of Turaev [21] show that each Yang–Baxter operator gives rise to an associated knot invariant. While these operators are therefore of high importance in knot theory, they are very difficult to construct; initially, the only source of solutions was the theory of quantum groups. However, Drinfeld [8] observed that a particular class of solutions provides a handier way to obtain classes of solutions by demanding that $R$ permutes the tensors $v_{i}\otimes v_{j}$ where $(v_{i})_{i\in I}$ is a basis of $V$. By doing so, one gets a discretized version of the Yang–Baxter equation, the set-theoretic Yang–Baxter equation, which consists of a set $X$ and a bijection $r:X^{2}\to X^{2}$ that satisfies the following equation in $\operatorname{Sym}_{X^{3}}$, the symmetric group on the set $X^{3}$:
@@
Indeed, set-theoretic solutions admit an analysis via algebraic methods, the first steps being made by Etingof, Schedler and Soloviev [9] and Gateva-Ivanova and Van den Bergh [11].
A homomorphism of cycle sets $f:(X,r)\to(Y,s)$ is a map $f:X\to Y$ such that $(f\times f)r=s(f\times f)$. An isomorphism of cycle sets is a bijective homomorphism.
Writing a set-theoretic solution $(X,r)$ as $r(x,y)=(\lambda_{x}(y),\rho_{y}(x))$ we call it
@@
    non-degenerate if the mappings $\lambda_{x},\rho_{x}$ are bijective for all $x\in X$,
@@
    involutive if $r^{2}=\operatorname{id}_{X^{2}}$.
By the size of a set-theoretic solution, we mean the size of $X$. From now on, when talking about set-theoretic solutions to the Yang–Baxter equation that are finite, involutive and non-degenerate we will most of the time simply write set-theoretic solutions or solutions.
@@
The cycle sets considered in this work will always be assumed to be of finite size and will be referred to as cycle sets, without the predicate non-degenerate. Cycle sets are in bijective correspondence with set-theoretic solutions to the Yang-Baxter equation. More precisely, given a cycle set $(X,\cdot)$ then $(X,r_{X})$ is a set-theoretic solution to the YBE, where $r_{X}:X^{2}\to X^{2}$ is given by
@@
Homomorphisms of cycle sets are defined in the obvious way. Consequently, this holds for all derived notions, such as isomorphisms. Note that isomorphisms of cycle sets translate to isomorphisms of set-theoretic solutions under the correspondence described above. As usual, if there is no danger of confusion, a cycle set $(X,\cdot)$ is simply referred to by its underlying set $X$.
Of particular importance for understanding general cycle sets is the class of indecomposable cycle sets. A cycle set $X$ is called indecomposable if there are no subsets $X_{1},X_{2}\neq\emptyset$ closed under the cycle set operation such that $X=X_{1}\cup X_{2}$. In a sense, they are elementary building blocks for cycle sets and turn out to be significantly more well-behaved. This can be seen from computations of Akgün, Mereb and Vendramin [1] that imply that the number of cycle sets of given cardinality grows in an aggressive manner, while the number of indecomposable cycle sets stays quite small. This indicates that there are strong obstructions for a cycle set to be indecomposable. Note that a cycle set $(X,\cdot)$ is indecomposable if and only if the corresponding set-theoretic solution $(X,r)$ is indecomposable in the sense that there is no partition $X=A\sqcup B$ ($A,B\neq\emptyset$) such that $r(A^{2})=A^{2}$ and $r(B^{2})=B^{2}$.
@@
In this article, we build on these results by providing a full classification of indecomposable cycle sets of size $p^{2}$, $p$ a prime. Using the correspondence between cycle sets and set-theoretic solutions, our final result, 5.1, reformulated in terms of set-theoretic solutions, will be as follows:
Theorem.
Each non-degenerate, indecomposable, involutive set-theoretic solution $(X,r)$ to the Yang–Baxter equation of size $p^{2}$ for some prime $p$ is isomorphic to one of the following solutions:
  1. (1)
    $X=\mathbb{Z}_{p^{2}}$, with $r(x,y)=(y+1,x-1)$.
  2. (2)
    $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, with
    $\chi_{0}(x)$ =$=\begin{cases}1&x=0\\ 0&x\neq 0\end{cases}.$
  3. (3)
    $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, with
@@ 1. Preliminaries
The group $A^{+}=(A,+)$ is the additive group of $A$ and $A^{\circ}=(A,\circ)$ is the multiplicative group of $A$. We will also denote the conjugation by $a$ in $A^{\circ}$ as ${}^{a}b=a\circ b\circ\overline{a}$ for all $a,b\in A$. For $n\in\mathbb{Z}$, $a\in A$, the $n$-th power of $a$ in $A^{+}$ is denoted $na$ and the $n$-th power in $A^{\circ}$ is denoted $a^{\circ n}$.
@@
A left ideal of $A$ is a subset $I$ of $A$ such that $I$ is a subgroup of $A^{+},$ and $\lambda_{a}(I)\subseteq I$ for all $a\in A$. Note that in particular, every characteristic subgroup of $A^{+}$ is a left ideal. If additionally, $I$ is a normal subgroup of $A^{\circ}$, one calls $I$ an ideal.
@@
In particular, if $a\in\operatorname{Fix}(A)$, then $\lambda_{a}(b)={}^{a}b$ for all $b\in A$ and thus $\lambda_{a}$ is a brace automorphism of $A$.
@@
Observe that $\lambda_{a}(b)=-a+a\circ b=-a+{}^{a}b\circ a=-a+{}^{a}b+\lambda_{{}^{a}b}(a)$. ∎
@@
We will not use this consequence but it is worth noting that this line of argumentation proves, more generally, the following statement for skew braces [12]:
Corollary 1.5.
Let $A$ be a finite skew brace and $B$ a subbrace such that $p=|A|/|B|$ is equal to the least prime dividing $|A|$. Then $B$ is an ideal of $A$.
@@
If $X$ is a set, we denote by $\operatorname{Sym}_{X}$ the symmetric group consisting of all permutations of $X$. If $n$ is a positive integer, we write $\operatorname{Sym}_{n}=\operatorname{Sym}_{\{1,\ldots,n\}}$.
@@
If there is more than one non-trivial system of imprimitivity, a theorem of Lucchini [16, Theorem 1] implies that $G\leqslant\operatorname{Sym}_{p}\times\operatorname{Sym}_{p}$. But if $G$ is a $p$-group, this forces $G$ to be isomorphic to a subgroup of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and therefore to be abelian. ∎
Let $G,H$ be groups such that $G$ acts on the set $X$. Then the wreath product is defined as the semidirect product $G\wr_{X}H=G\ltimes H^{X}$ where $G$ acts on $H^{X}$ by $g\cdot(h_{x})_{x\in X}=(h_{g^{-1}\cdot x})_{x\in X}$. If the action of $G$ on $X$ is clear, we will generally suppress the subscript-$X$ and write $G\wr H$.
@@
Let $G\leqslant\operatorname{Sym}_{\mathbb{Z}_{p}\times\mathbb{Z}_{p}}$ be a transitive solvable group such that the sets $\{a\}\times\mathbb{Z}_{p}$, $a\in\mathbb{Z}_{p}$, form a system of imprimitivity, then $G$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. If moreover $G$ is a $p$-group, then $G$ is conjugate to a subgroup of $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$.
@@
As a special case of [13, II, Satz 3.2], we know that every solvable transitive subgroup of $\operatorname{Sym}_{\mathbb{Z}_{p}}$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)$. So we conclude that $G$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. As $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$ is a $p$-Sylow subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$ also the last part of the statement follows. ∎
@@
with $\beta,\gamma_{1},\ldots,\gamma_{p}\in\mathbb{Z}_{p}$.
On any brace $A$ we have a canonical cycle set structure $(A,\cdot)$ where $a\cdot b=\lambda_{a}^{-1}(b)$. A subset $X\subseteq A$ is a cycle base of $A$ if it is a sub-cycle set of $(A,\cdot)$ and also $X$ generates the groups $A^{+}$ and $A^{\circ}$. A cycle base is called transitive if $X$ forms an orbit under the $\lambda$-action of $A$, or equivalently if $X$ is an indecomposable cycle set.
@@ 2. Indecomposable retractable cycle sets of size $p^{2}$
For the reader’s convenience, we first recall a special case of the construction of indecomposable cycle sets of multipermutation level $2$ by Jedlička and Pilitowska [14, Proposition 5.1, Proposition 5.7].
@@
    a subgroup $H\leqslant\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$ such that $c_{i}-c_{i+m}\in H$ for all $i\in\mathbb{Z}$,
@@
It follows from 1.9 that $|\operatorname{Ret}(X)|\in\{1,p\}$. If $|\operatorname{Ret}(X)|=1$ then $X$ clearly has finite multipermutation level and is isomorphic to the given cyclic cycle set on $\mathbb{Z}_{p^{2}}$. If $|\operatorname{Ret}(X)|=p$, then we know that $\operatorname{Ret}(X)$ is isomorphic to the cycle set on $\mathbb{Z}_{p}$ with $x\cdot y=y+1$ [9, Theorem 2.13]. In particular, $|\operatorname{Ret}(\operatorname{Ret}(X))|=1$ and thus $X$ has multipermutation level 2. The statement now follows directly from 2.2. ∎
@@ 3. Indecomposable irretractable cycle sets of size $p^{2}$ The $p$-group case
In this section, we will solve the following problem:
Problem 3.1.
Let $p$ be a prime. Find all isomorphism classes of indecomposable irretractable cycle sets $(X,\cdot)$ of size $p^{2}$ such that $\mathcal{G}(X)$ is a $p$-group.
We first construct these solutions in Section 3.1. In Section 3.2 we then study when such solutions are isomorphic and moreover we determine their automorphism groups.
@@ 3.1. Constructing the solutions
We first observe the following:
@@
As $X$ is irretractable, it follows from [2, Lemma 2.1.] that $\operatorname{Soc}(\mathcal{G}(X))=0$. From 1.6 we then find that $\mathcal{G}(X)^{\circ}$ is not abelian. ∎
@@
By 1.7 and 3.2 we have a unique system of imprimitivity, for $x\in X$ we denote by $\mathcal{B}_{x}$ the block containing $x$. We denote by $\mathcal{A}=\mathcal{G}\cap\mathbb{Z}_{p}^{p}$ the abelian subgroup of $\mathcal{G}^{\circ}$ which fixes the blocks setwise. Note that $[\mathcal{G}^{\circ}:\mathcal{A}]=p$.
@@
Suppose that $0\neq f\in\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}$, without loss of generality we may assume that $f^{\circ p}=0$. By 1.1 and the abelianity of $\mathcal{A}$, we see that $\lambda_{f}(g)={}^{f}g=g$ for all $g\in\mathcal{A}$. Thus $\lambda_{f}$ fixes all elements in $\mathcal{A}$. Moreover, by 1.2, $\operatorname{Fix}(\mathcal{G})\cap Z(\mathcal{G}^{\circ})\subseteq\operatorname{Soc}(\mathcal{G}^{\circ})=0$, so $f\not\in Z(\mathcal{G}^{\circ})$. Hence, knowing also that $|\mathcal{G}|/|\mathcal{A}|=p$, we see that $\mathcal{A}=\{g\in\mathcal{G}\mid{}^{f}g=g\}$. As $\lambda_{f}$ is an automorphism, $\mathcal{A}$ is a subbrace of index $p$ in $\mathcal{G}$. By 1.4, $\mathcal{A}$ is an ideal of $\mathcal{G}$.
@@
In particular, this implies that $p\gamma=0$. If $\gamma=0$, then $\lambda_{f}(x)=x$ for all $x\in X$ and therefore $\lambda_{f}=\operatorname{id}_{\mathcal{G}}$, which would imply that $f\in\operatorname{Soc}(\mathcal{G})$. It therefore follows that $\gamma\neq 0$, which in turn implies that $\lambda_{f}$ has no fixed points on $X$. In particular, for each $a\in\mathcal{A}$ and $x\in X$, there exists some $n\in\mathbb{N}$ such that $\lambda_{a}(x)=\lambda_{f}^{n}(x)$, hence $a*x=f^{\circ n}*x=n\gamma$.
Now consider $I$ be the subgroup of $\mathcal{G}^{+}$ generated by $\{a*g\mid a\in\mathcal{A},g\in\mathcal{G}\}$, which is an ideal by [18, Corollary after Proposition 6]. Because $X$ generates $\mathcal{G}^{+}$ and * is left distributive, $I$ is the subgroup of $\mathcal{G}^{+}$ generated by $\{a*x\mid a\in\mathcal{A},x\in X\}$. By the previous discussion, $I=\{0,\gamma,\ldots,(p-1)\gamma\}$. As $|I|=p$, we find that $I\subseteq\operatorname{Fix}(\mathcal{G})$. Since $I$ is a minimal normal subgroup of the nilpotent group $\mathcal{G}^{\circ}$, we also find that $I\subseteq Z(\mathcal{G}^{\circ})$, hence 1.2 yields a contradiction with the assumption that $\operatorname{Soc}(\mathcal{G})=0$. ∎
Now that we know that $\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}=0$, it follows from 1.6 that $|\operatorname{Fix}(\mathcal{G})|=p$. Also, we have a semidirect product $\mathcal{G}^{\circ}=\mathcal{A}\rtimes\operatorname{Fix}(\mathcal{G})^{\circ}$. We use this for the following construction: for all $g\in\mathcal{G}$, we have an equality $g\circ\operatorname{Fix}(\mathcal{G})=g+\lambda_{g}(\operatorname{Fix}(\mathcal{G}))=g+\operatorname{Fix}(\mathcal{G})$. Therefore, $\mathcal{A}$ forms a system of representatives for $\mathcal{G}^{\circ}/\operatorname{Fix}(\mathcal{G})^{\circ}$ and $\mathcal{G}^{+}/\operatorname{Fix}(\mathcal{G})^{+}$. We keep the multiplication on $\mathcal{A}$ as it is, but as $\mathcal{A}$ is not necessarily closed under $+$, we define $g\oplus h$ as the unique element in $\mathcal{A}\cap(g+h+\operatorname{Fix}(\mathcal{G}))$.
@@
$\mathcal{A}$ is a brace under $\circ$ and $\oplus$.
@@
Suppose that $\tau:X\to\tilde{X}$ is injective, then $\mathcal{G}(\tilde{X})^{\circ}$ is isomorphic to $\mathcal{A}$ and each element of $\tilde{X}$ acts differently on $\tilde{X}$ by the $\tilde{\lambda}$-action, so $\tilde{X}$ is irretractable. Hence, this contradicts 3.2. Therefore, $|\tilde{X}|=p$. ∎
@@
By 3.5, for all $x\in X$, we have $x+\operatorname{Fix}(\mathcal{G})\subseteq X$. As $\mathcal{G}^{\circ}$ leaves $\operatorname{Fix}(\mathcal{G})$ invariant under the $\lambda$-action, we deduce that the cosets $x+\operatorname{Fix}(\mathcal{G})$ form a non-trivial system of imprimitivity for the $\lambda$-action of $\mathcal{G}^{\circ}$ on $X$. The uniqueness of such a system (1.7) implies that $\mathcal{B}_{x}=x+\operatorname{Fix}(\mathcal{G})$.

### v3

@@ Abstract.
This article focuses on indecomposable involutive non-degenerate set-theoretical solutions to the Yang–Baxter equation. More specifically, we give a full classification of those solutions which are of size $p^{2}$, for $p$ a prime. We do this through a thorough analysis of their associated permutation braces and using the language of cycle sets.
Keywords: brace, cycle set, Yang–Baxter equation
@@ Introduction
Outside of physics, solutions to this equation have been recognized to be of importance in low-dimensional topology, as results of Turaev [20] show that each Yang–Baxter operator gives rise to an associated knot invariant. While these operators are therefore of high importance in knot theory, they are very difficult to construct; initially, the only source of solutions was the theory of quantum groups. However, Drinfeld [8] observed that a particular class of solutions provides a handier way to obtain classes of solutions by demanding that $R$ permutes the tensors $v_{i}\otimes v_{j}$ where $(v_{i})_{i\in I}$ is a basis of $V$. By doing so, one gets a discretized version of the Yang–Baxter equation, the set-theoretical Yang–Baxter equation, which consists of a set $X$ and a bijection $r:X^{2}\to X^{2}$ that satisfies the following equation in $\operatorname{Sym}_{X^{3}}$, the symmetric group on the set $X^{3}$:
@@
Indeed, set-theoretical solutions admit an analysis via algebraic methods, the first steps being made by Etingof, Schedler and Soloviev [9] and Gateva-Ivanova and Van den Bergh [11].
A homomorphism of solutions $f:(X,r)\to(Y,s)$ is a map $f:X\to Y$ such that $(f\times f)r=s(f\times f)$. An isomorphism of solutions is a bijective homomorphism.
Writing a set-theoretical solution $(X,r)$ as $r(x,y)=(\lambda_{x}(y),\rho_{y}(x))$ we call it
@@
    non-degenerate, if the mappings $\lambda_{x},\rho_{x}$ are bijective for all $x\in X$,
@@
    involutive, if $r^{2}=\operatorname{id}_{X^{2}}$.
By the size of a set-theoretical solution, we mean the size of $X$.
Convention.
From now on, we simply write set-theoretical solutions or solutions when talking about set-theoretical solutions to the Yang–Baxter equation that are finite, involutive and non-degenerate.
@@
Convention.
The cycle sets considered in this work will always be assumed to be of finite size and will be referred to as cycle sets, without the predicate non-degenerate.
Cycle sets are in bijective correspondence with set-theoretical solutions to the Yang–Baxter equation. More precisely, given a cycle set $(X,\cdot)$ then $(X,r_{X})$ is a set-theoretical solution to the YBE, where $r_{X}:X^{2}\to X^{2}$ is given by
@@
Homomorphisms of cycle sets are defined in the obvious way. Consequently, this holds for all derived notions, such as isomorphisms. Note that isomorphisms of cycle sets translate to isomorphisms of set-theoretical solutions under the correspondence described above. As usual, if there is no danger of confusion, a cycle set $(X,\cdot)$ is simply referred to by its underlying set $X$.
Of particular importance for understanding general cycle sets is the class of indecomposable cycle sets. A cycle set $X$ is called indecomposable if there are no subsets $X_{1},X_{2}\neq\emptyset$ closed under the cycle set operation such that $X=X_{1}\cup X_{2}$. In a sense, they are elementary building blocks for cycle sets and turn out to be significantly more well-behaved. This can be seen from computations of Akgün, Mereb and Vendramin [1] that imply that the number of cycle sets of given cardinality grows in an aggressive manner, while the number of indecomposable cycle sets stays quite small. This indicates that there are strong obstructions for a cycle set to be indecomposable. Note that a cycle set $(X,\cdot)$ is indecomposable if and only if the corresponding set-theoretical solution $(X,r)$ is indecomposable in the sense that there is no partition $X=A\sqcup B$ ($A,B\neq\emptyset$) such that $r(A^{2})=A^{2}$ and $r(B^{2})=B^{2}$.
@@
In this article, we build on these results by providing in 5.1 a full classification of indecomposable cycle sets of size $p^{2}$, $p$ a prime. In 5.2 we reformulate this classification in terms of set-theoretical solutions.
@@ 1. Preliminaries
The group $A^{+}=(A,+)$ is the additive group of $A$ and $A^{\circ}=(A,\circ)$ is the multiplicative group of $A$. We will also denote the conjugation by $a$ in $A^{\circ}$ as ${}^{a}b=a\circ b\circ\overline{a}$, for all $a,b\in A$. For $n\in\mathbb{Z}$, $a\in A$, the $n$-th power of $a$ in $A^{+}$ is denoted $na$ and the $n$-th power in $A^{\circ}$ is denoted $a^{\circ n}$.
@@
A left ideal of $A$ is a subset $I$ of $A$ such that $I$ is a subgroup of $A^{+},$ and $\lambda_{a}(I)\subseteq I$, for all $a\in A$. Note that, in particular, every characteristic subgroup of $A^{+}$ is a left ideal. If additionally, $I$ is a normal subgroup of $A^{\circ}$, one calls $I$ an ideal of $A$.
@@
In particular, if $a\in\operatorname{Fix}(A)$, then $\lambda_{a}(b)={}^{a}b$, for all $b\in A$ and thus $\lambda_{a}$ is a brace automorphism of $A$.
@@
Observe that if $a,b\in A$, then
$\lambda_{a}(b)=-a+a\circ b=-a+{}^{a}b\circ a=-a+{}^{a}b+\lambda_{{}^{a}b}(a).\qed$
@@
If $X$ is a set, we denote by $\operatorname{Sym}_{X}$ the symmetric group consisting of all permutations of $X$. If $n$ is a positive integer, we write $\operatorname{Sym}_{n}$ to indicate $\operatorname{Sym}_{\{1,\ldots,n\}}$.
@@
If there is more than one non-trivial system of imprimitivity, a theorem of Lucchini [15, Theorem 1] implies that $G\leqslant\operatorname{Sym}_{p}\times\operatorname{Sym}_{p}$. But if $G$ is a $p$-group, this forces $G$ to be isomorphic to a subgroup of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and therefore to be abelian. ∎
Let $G,H$ be groups such that $G$ acts on a set $X$. The wreath product is the semidirect product $G\wr_{X}H=G\ltimes H^{X}$, where $H^{X}=\{(h_{x})_{x\in X}\mid h_{x}\in H\text{ for all }x\in X\}$ is the iterated direct product of $H$ with itself indexed by $X$ and $G$ acts on $H^{X}$ by $g\cdot(h_{x})_{x\in X}=(h_{g^{-1}\cdot x})_{x\in X}$. If the action of $G$ on $X$ is clear, we will generally suppress the subscript-$X$ and write $G\wr H$.
@@
Let $G\leqslant\operatorname{Sym}_{\mathbb{Z}_{p}\times\mathbb{Z}_{p}}$ be a transitive solvable group such that the sets $\{a\}\times\mathbb{Z}_{p}$, $a\in\mathbb{Z}_{p}$, form a system of imprimitivity, then $G$ is conjugated to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. If moreover $G$ is a $p$-group, then $G$ is conjugated to a subgroup of $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$.
@@
As a special case of [12, Chapter II, Satz 3.2], we know that every solvable transitive subgroup of $\operatorname{Sym}_{\mathbb{Z}_{p}}$ is conjugated to a subgroup of $\mathrm{AGL}(1,p)$. So we conclude that $G$ is conjugated to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. As $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$ is a $p$-Sylow subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$ also the last part of the statement follows. ∎
@@
with $\beta,\gamma_{0},\ldots,\gamma_{p-1}\in\mathbb{Z}_{p}$.
On any brace $A$ we have a canonical cycle set structure $(A,\cdot)$ where $a\cdot b=\lambda_{a}^{-1}(b)$, for all $a,b\in A$. A subset $X\subseteq A$ is a cycle base of $A$ if it is a sub-cycle set of $(A,\cdot)$ and also $X$ generates the groups $A^{+}$ and $A^{\circ}$. A cycle base is called transitive if $X$ forms an orbit under the $\lambda$-action of $A$, or equivalently if $X$ is an indecomposable cycle set.
@@ 2. Indecomposable retractable cycle sets of size $p^{2}$
For the reader’s convenience, we first recall a special case of the construction of indecomposable cycle sets of multipermutation level $2$ by Jedlička and Pilitowska [13, Proposition 5.1, Proposition 5.7].
@@
    a subgroup $H\leqslant\bigoplus_{i\in\mathbb{Z}}\mathbb{Z}$ such that $c_{i}-c_{i+m}\in H$, for all $i\in\mathbb{Z}$,
@@
It follows from 1.8 that $|\operatorname{Ret}(X)|\in\{1,p\}$. If $|\operatorname{Ret}(X)|=1$ then $X$ clearly has finite multipermutation level and is isomorphic to the given cyclic cycle set on $\mathbb{Z}_{p^{2}}$. If $|\operatorname{Ret}(X)|=p$, then we know that $\operatorname{Ret}(X)$ is isomorphic to the cycle set on $\mathbb{Z}_{p}$ with $x\cdot y=y+1$ [9, Theorem 2.13]. In particular, $|\operatorname{Ret}(\operatorname{Ret}(X))|=1$ and thus $X$ has multipermutation level 2. The statement now follows directly from 2.2. ∎
@@ 3. Indecomposable irretractable cycle sets of size $p^{2}$ The $p$-group case
The aim of this section is to find all isomorphism classes of indecomposable irretractable cycle sets $(X,\cdot)$ of size $p^{2}$ such that $\mathcal{G}(X)$ is a $p$-group, for $p$ a prime number.
We first construct these cycle sets in Section 3.1. In Section 3.2 we then study when such cycle sets are isomorphic and moreover we determine their automorphism groups.
@@ 3.1. Constructing the solutions
We first introduce the following result.
@@
As $X$ is irretractable, it follows from [2, Lemma 2.1] that $\operatorname{Soc}(\mathcal{G}(X))=0$. From 1.5 we then find that $\mathcal{G}(X)^{\circ}$ is not abelian. ∎
@@
By 1.6 and 3.1 we have a unique system of imprimitivity, for $x\in X$ we denote by $\mathcal{B}_{x}$ the block containing $x$. We denote by $\mathcal{A}=\mathcal{G}\cap\mathbb{Z}_{p}^{p}$ the abelian subgroup of $\mathcal{G}^{\circ}$ which fixes the blocks setwise. Note that $[\mathcal{G}^{\circ}:\mathcal{A}]=p$.
@@
Suppose that $0\neq f\in\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}$, without loss of generality we may assume that $f^{\circ p}=0$. By 1.1 and the abelianity of $\mathcal{A}$, we see that $\lambda_{f}(g)={}^{f}g=g$, for all $g\in\mathcal{A}$. Thus $\lambda_{f}$ fixes all elements in $\mathcal{A}$. Moreover, by 1.2, $\operatorname{Fix}(\mathcal{G})\cap Z(\mathcal{G}^{\circ})\subseteq\operatorname{Soc}(\mathcal{G}^{\circ})=0$, so $f\not\in Z(\mathcal{G}^{\circ})$. Hence, knowing also that $|\mathcal{G}|/|\mathcal{A}|=p$, we see that $\mathcal{A}=\{g\in\mathcal{G}\mid{}^{f}g=g\}$. As $\lambda_{f}$ is an automorphism, $\mathcal{A}$ is a subbrace of index $p$ in $\mathcal{G}$. By 1.4, $\mathcal{A}$ is an ideal of $\mathcal{G}$.
@@
In particular, this implies that $p\gamma=0$. If $\gamma=0$, then $\lambda_{f}(x)=x$, for all $x\in X$ and therefore $\lambda_{f}=\operatorname{id}_{\mathcal{G}}$, which would imply that $f\in\operatorname{Soc}(\mathcal{G})$. It therefore follows that $\gamma\neq 0$, which in turn implies that $\lambda_{f}$ has no fixed points on $X$. In particular, for each $a\in\mathcal{A}$ and $x\in X$, there exists some $n\in\mathbb{N}$ such that $\lambda_{a}(x)=\lambda_{f}^{n}(x)$, hence $a*x=f^{\circ n}*x=n\gamma$.
Now consider $I$ be the subgroup of $\mathcal{G}^{+}$ generated by $\{a*g\mid a\in\mathcal{A},g\in\mathcal{G}\}$, which is an ideal by [17, Corollary after Proposition 6]. Because $X$ generates $\mathcal{G}^{+}$ and * is left distributive, $I$ is the subgroup of $\mathcal{G}^{+}$ generated by $\{a*x\mid a\in\mathcal{A},x\in X\}$. By the previous discussion, $I=\{0,\gamma,\ldots,(p-1)\gamma\}$. As $|I|=p$, we find that $I\subseteq\operatorname{Fix}(\mathcal{G})$. Since $I$ is a minimal normal subgroup of the nilpotent group $\mathcal{G}^{\circ}$, we also find that $I\subseteq Z(\mathcal{G}^{\circ})$, hence 1.2 yields a contradiction with the assumption that $\operatorname{Soc}(\mathcal{G})=0$. ∎
Now that we know that $\operatorname{Fix}(\mathcal{G})\cap\mathcal{A}=0$, it follows from 1.5 that $|\operatorname{Fix}(\mathcal{G})|=p$. Also, we have a semidirect product $\mathcal{G}^{\circ}=\mathcal{A}\rtimes\operatorname{Fix}(\mathcal{G})^{\circ}$. We use this for the following construction: for all $g\in\mathcal{G}$, we have an equality $g\circ\operatorname{Fix}(\mathcal{G})=g+\lambda_{g}(\operatorname{Fix}(\mathcal{G}))=g+\operatorname{Fix}(\mathcal{G})$. Therefore, $\mathcal{A}$ forms a system of representatives for $\mathcal{G}^{\circ}/\operatorname{Fix}(\mathcal{G})^{\circ}$ and $\mathcal{G}^{+}/\operatorname{Fix}(\mathcal{G})^{+}$. We keep the multiplication on $\mathcal{A}$ as it is, but as $\mathcal{A}$ is not necessarily closed under $+$, we define $g\oplus h$ as the unique element in $\mathcal{A}\cap(g+h+\operatorname{Fix}(\mathcal{G}))$.
@@
The structure $(\mathcal{A},\oplus,\circ)$ is a brace.
@@
Suppose that $\tau:X\to\tilde{X}$ is injective, then $\mathcal{G}(\tilde{X})^{\circ}$ is isomorphic to $\mathcal{A}$ and each element of $\tilde{X}$ acts differently on $\tilde{X}$ by the $\tilde{\lambda}$-action, so $\tilde{X}$ is irretractable. Hence, this contradicts 3.1. Therefore, $|\tilde{X}|=p$. ∎
@@
By 3.4, for all $x\in X$, we have $x+\operatorname{Fix}(\mathcal{G})\subseteq X$. As $\mathcal{G}^{\circ}$ leaves $\operatorname{Fix}(\mathcal{G})$ invariant under the $\lambda$-action, we deduce that the cosets $x+\operatorname{Fix}(\mathcal{G})$ form a non-trivial system of imprimitivity for the $\lambda$-action of $\mathcal{G}^{\circ}$ on $X$. The uniqueness of such a system (1.6) implies that $\mathcal{B}_{x}=x+\operatorname{Fix}(\mathcal{G})$.